given 3x- 4y = 7 and x+ cy = 13 for what value of c will this system of equations NOT have a solution?
a -4 b -4/3 c -3/4 d 3/4 e 4/3

Answer by edjones(8007)

3x- 4y = 7
x+ cy = 13
.
-4y=-3x+7
y=3x/4 -7/4
.
cy=13-x
y=13/c - x/c
.
There would be no answer if the equations had the same slope:
3x/4=-x/c Let the slope of the 2nd equal the slope of the 1st.
c/4=-x/3x Multiply each side by c/3x
c=-4/3
.
Ed